# From Interaction to Impact: Towards Safer AI Agents Through Understanding and Evaluating UI Operation Impacts

### v1

@@ Abstract.
With advances in generative AI, there is increasing work towards creating autonomous agents that can manage daily tasks by operating user interfaces (UIs). While prior research has studied the mechanics of how AI agents might navigate UIs and understand UI structure, the effects of agents and their autonomous actions—particularly those that may be risky or irreversible—remain under-explored. In this work, we investigate the real-world impacts and consequences of UI actions by AI agents. We began by developing a taxonomy of the impacts of UI actions through a series of workshops with domain experts. Following this, we conducted a data synthesis study to gather realistic UI screen traces and action data that users perceive as impactful. We then used our impact categories to annotate our collected data and data repurposed from existing UI navigation datasets. Our quantitative evaluations of different large language models (LLMs) and variants demonstrate how well different LLMs can understand the impacts of UI actions that might be taken by an agent. We show that our taxonomy enhances the reasoning capabilities of these LLMs for understanding the impacts of UI actions, but our findings also reveal significant gaps in their ability to reliably classify more nuanced or complex categories of impact.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing HCI theory, concepts and modelsccs: Computing methodologies Machine learning
Figure 1. We studied and modeled the impacts of UI operations through (a) a workshop study that iterated a taxonomy categorizing the impacts, (b) a data synthesis study to collect UI actions with impacts, and (c) evaluation of LLM implementations on their capabilities of understanding impacts. We contribute the taxonomy and the evaluation findings for future AI agents to better understand the consequences of UI actions.
@@ 1. Introduction
While prior research has explored the automation of both physical tasks and digital tasks on user interfaces (Huang et al., 2022; Wang et al., 2023b; Imran et al., 2024; Zhou et al., 2023; Wang et al., 2024; Gur et al., 2023), much of the focus in digital automation has been on enabling AI to understand UI screens, plan specific actions, and navigate through user interfaces (Wen et al., 2023a; Taeb et al., 2023; Wen et al., 2023b; Burns et al., 2022b; Zhang et al., 2021). However, one critical aspect remains underexplored: understanding the consequences of agents performing UI actions, especially those may be risky or irreversible, even when directed by high-level user instructions. In today’s digital age, where interactions with the world are increasingly mediated through screens, it is crucial for AI agents to anticipate and reason about the potential impacts and consequences of their actions on users and their environments.
@@
This paper presents a significant step towards addressing this gap by first constructing a comprehensive taxonomy of UI action impacts. Through a series of workshops with 12 domain experts in LLMs, UI understanding, and AI safety, we iteratively developed this taxonomy, identifying and categorizing UI actions with potentially significant effects from different domains like impact on the user themselves, impact on other users, the reversibility of this action, etc. In this work, we focus on the impacts caused by attempting to carry out human-initiated directions. In our work we exclude potential passive impacts, such as recommendation algorithms that change what users see based on their browsing activities, as such passive impacts are subtle and hard to quantify.
@@
After we established the taxonomy, we conducted a data synthesis study to collect realistic UI actions likely to have impacts by recording UI screen traces and users’ intended actions. We show how existing UI datasets like MoTIF (Burns et al., 2022a) and AndroidControl (Li et al., 2024) are mostly “harmless” browsing and searching tasks, while our generated UI action traces contain tasks with potential impacts (e.g., changing account information, sending a message, etc.)
@@
We collected and labeled 250 UI action traces using our taxonomy via crowdsourcing. Together with 1319 pieces of data sampled from MoTIF and AndroidControl, we evaluated whether LLMs can predict the impact level in comparison to a human’s perceived judgment and accurately classify categories as outlined in our taxonomy. We differentiated the impact levels based on the degree of user intervention required, ranging from users comfortable with an action being automated to those necessitating direct user control. We found that knowledge of our taxonomy could increase an LLM’s ability to predict the overall impact level. However, while existing LLMs demonstrate a moderate understanding on some categories in the taxonomy and predict reasonable results, they still fail to understand all of the nuanced aspects of a UI action’s potential impact. We also present additional findings on how the LLMs we tested overestimated impact and the differences in perceived impact.
@@
  • The development of a comprehensive taxonomy categorizing the impact of UI actions,
  • A collection of human-synthesized UI action traces that include actions with potentially negative impacts, annotated together with data sampled from two other existing datasets,
  • An evaluation and discussion of how a representative set of current LLMs assess the impact of UI actions, showing a significant gap in their ability to reliably understand the complexity of the impact of UI actions.
@@ 2.1. AI Agents and Safety
LLM-based techniques are moving from limited chat bots to agents (Park et al., 2023; Wang et al., 2023a; Liu et al., 2023b, c; Ahn et al., 2022; Driess et al., 2023a) which have the capability to interact with both physical and digital environments. The functionality of these agents also varies based on their initiatives and agencies. Extensive works have explored using LLMs as assistants for creativity tasks including writing (Lee et al., 2022; Chakrabarty et al., 2022), coding (Wu et al., 2024; Agarwal et al., 2024), and content generation (Chang et al., 2023; Lin et al., 2023; Kondratyuk et al., 2023). In the meantime, a number of research efforts have started to explore autonomous AI agents with stronger initiatives and agencies. These agents have been employed in tasks including autonomous decision-making in dynamic environments (Wu et al., 2023; Hong et al., 2023; Liu et al., 2023d), adaptive problem-solving (Lu et al., 2023; do Nascimento et al., 2023), conducting research (Dai et al., 2023b), manipulating digital interfaces (Wen et al., 2023a; Ruan et al., 2023a), and even interacting with the real world via robotics and IoT systems (Xiao and Wang, 2023; Yang et al., 2023; Cui et al., 2023). These advancements brought us back to the discussion which emerged in earlier years of AI and HCI research, i.e., the dynamics between users and system initiatives (Horvitz, 1999; Hartmann, 2009; Amershi et al., 2019). More recently, as LLM-based agents’ capabilities have been advanced in different tasks, it is more crucial to address the problem of identifying and reacting to consequences or risks of agent actions, and introduce human confirmation or intervention when necessary. Works focusing on the trustworthiness and safety (Sun et al., 2024; et al., 2022; Glaese et al., 2022) of such agents include ToolEmu (Ruan et al., 2023b) which used an language model emulated sandbox to identify the risks of agents, and TrustAgent (Hua et al., 2024) which used three strategic planning to inject safety knowledge, generate safe plans, and inspect post-planning.
Our work differs from prior research on safer AI agents by focusing on what would happen after AI agents perform UI actions on smartphone screens as people rely heavily on them for interacting with the world.
@@ 2.2. UI Understanding, Planning, and Navigation
Prior work has largely explored UI understanding, planning, and navigation on UI screens, which provided a substantial grounding for this work. Prior to the era of LLMs, earlier works in this area studied simpler web and mobile UI screens (Burns et al., 2022a; Gur et al., 2018; Li et al., 2020; Liu et al., 2018; Shi et al., 2017). Zhang et al. proposed Screen Recognition (Zhang et al., 2021), which predicts bounding boxes, labels, text content, and clickability of UI elements from screenshot pixels. More recently, both LLM (et al., 2023; Driess et al., 2023b; Gu and Dao, 2023; Huang et al., 2023; Jiang et al., 2023a; OpenAI, 2023; Touvron et al., 2023) and multimodal LLM (MLLM) (Dai et al., 2023a; Li et al., 2023b, a; Liu et al., 2023a; McKinzie et al., 2024; Sun et al., 2023; Ye et al., 2023; Zhu et al., 2023) research have advanced the capabilities of UI understanding tasks, including ILuvUI (Jiang et al., 2023b) and Spotlight (Li and Li, 2022) which focus on single-screen UI tasks like screen summarization and widget interaction via GPT-generated data. Ferret-UI (You et al., 2023) introduces strong referring, grounding, and reasoning capabilities to the UI domain.
@@ 3. Building a Taxonomy of UI Impacts
To better understand and categorize the consequences of an autonomous agent taking actions on an app, we created a preliminary taxonomy of the impacts of UI actions. We then conducted a set of workshop studies to elicit ideas on defining and categorizing UI action effects with 12 participants with expertise in LLMs, AI safety, and UI understanding. Over 4 workshop sessions, participants iterated our initial taxonomy into a refined version.
@@ 3.1. Author-designed Preliminary Taxonomy
Prior to the workshop sessions, the research team met and developed an initial version of a taxonomy designed to categorize the potential impacts of UI actions. The team achieved the initial taxonomy through a pilot workshop study to look at screenshots from existing datasets and brainstorm factors that would affect their decisions on why or why not an action has impact. To make it possible to model and categorize the broad concept of “impact,” the research team grounded the definition of “impact” to be any real world consequences that result directly from active user-initiated actions (e.g., clicking a confirm button or taking a screenshot). Any subtle or passive impact caused by underlying technologies behind the scene is not within our research scope. For example, scrolling through social media might cause future recommended feeds to change, but we are excluding these kind of impacts which are not caused actively by users. We hope to build a taxonomy to categorize the user-initiated active impact as a first step to understand the consequences of a UI action. We structured this initial taxonomy around the following key domains:
@@ 3.2.1. Participants
We recruited 12 domain experts from a large technology company through internal message boards and snowball sampling. The recruitment process began by contacting researchers with expertise in AI, and subsequently extended to those with specialized knowledge in AI safety and UI understanding.
@@ 3.3. Taxonomy Iteration
In this section, we present the findings from our workshop sessions, highlighting the depth and breadth of our discussions on UI action effects. The taxonomy we developed is intended to cover a wide range of scenarios that extend beyond just determining whether an AI agent should seek human confirmation or intervention. Instead, it broadly considers the various consequences that may arise after a UI action is executed, ensuring its applicability to diverse use cases.
@@ 4.1.2. Apparatus
Figure 4. An annotated example of a monetary transaction. Each of the category is from the taxonomy.
@@ 4.1.3. Procedure
4.2. Annotation
After collecting the UI action traces, which included both UI action traces and descriptions of the intended actions, we annotated them using our taxonomy through crowdsourcing. The resulting labeled dataset enabled us to evaluate the ability of AI agents to reason about the consequences of UI actions.
Additionally, we asked the annotators to to rate the impact of each action as minimal, moderate, or significant. We carefully defined the different levels of impact in the training process. We asked the annotators to rate an action as minimum impact if they felt like this action can be safely executed automatically without their confirmation. They were asked to rate it as moderate impact if they felt like they have some concerns over this action and would like to confirm or receive summary of the action before execution. The summary of the action refers to situations when actions contain rich information on the screen that require summarization and confirmation (e.g., shopping cart items before purchasing). Lastly, they were asked to rate the action as significant impact if they felt like they would not allow this action to be executed automatically and would like to intervene themselves. A typical action with significant impact is deleting an account that cannot be restored, where annotators felt that they themselves should be executing it if needed.
Each task was annotated by two annotators. In cases of disagreement, a third annotator was brought in to resolve the discrepancy. The third annotator was responsible for not only providing an additional opinion on the classification and rating questions, but also addressing the cases when the justifications were not similar.
5. Evaluation: Can LLMs Understand Impact
In this section, we aim to answer this research question: How do current LLMs understand the impact of UI actions? We summarize the collected data which potentially has impacts and quantitatively compare two existing datasets. We then assessed performance of five state-of-the-art LLMs with four different LLM prompting strategies assessing their abilities to determine UI action impact based on the taxonomy and to classify specific categories in the taxonomy.
5.1. Data Summary
We collected 250 unique UI action traces from various domains, including e-commerce, social media, financial services, productivity tools, travel and transportation, smart-homes, health, lifestyle, and others. After cleaning identical or highly similar screenshots that were accidentally captured in the collection phase, there are on average 5.3 screens per action trace. In total we collected 1328 UI screens. In the data annotation phase, annotators marked 41 action traces as incomplete and therefore leaving 209 data pieces with annotations. Within the collected data, annotators deemed 23.9% of them to have significant impact and 49.3% of them to have moderate impact, while the remaining 26.8% have minimum impact.

### v2

# From Interaction to Impact: Towards Safer AI Agents Through Understanding and Evaluating Mobile UI Operation Impacts

@@ Abstract.
With advances in generative AI, there is increasing work towards creating autonomous agents that can manage daily tasks by operating user interfaces (UIs). While prior research has studied the mechanics of how AI agents might navigate UIs and understand UI structure, the effects of agents and their autonomous actions—particularly those that may be risky or irreversible—remain under-explored. In this work, we investigate the real-world impacts and consequences of mobile UI actions taken by AI agents. We began by developing a taxonomy of the impacts of mobile UI actions through a series of workshops with domain experts. Following this, we conducted a data synthesis study to gather realistic mobile UI screen traces and action data that users perceive as impactful. We then used our impact categories to annotate our collected data and data repurposed from existing mobile UI navigation datasets. Our quantitative evaluations of different large language models (LLMs) and variants demonstrate how well different LLMs can understand the impacts of mobile UI actions that might be taken by an agent. We show that our taxonomy enhances the reasoning capabilities of these LLMs for understanding the impacts of mobile UI actions, but our findings also reveal significant gaps in their ability to reliably classify more nuanced or complex categories of impact.
@@
journalyear: 2025copyright: ccconference: 30th International Conference on Intelligent User Interfaces; March 24–27, 2025; Cagliari, Italybooktitle: 30th International Conference on Intelligent User Interfaces (IUI ’25), March 24–27, 2025, Cagliari, Italydoi: 10.1145/3708359.3712153isbn: 979-8-4007-1306-4/25/03ccs: Human-centered computing HCI theory, concepts and modelsccs: Computing methodologies Machine learning
Figure 1. We studied and modeled the impacts of UI operations on mobile interfaces through (a) a workshop study that iterated a taxonomy categorizing the impacts, (b) a data synthesis study to collect UI actions with impacts, and (c) evaluation of LLM implementations on their capabilities of understanding impacts. We contribute the taxonomy and the evaluation findings for future AI agents to better understand the consequences of UI actions.
@@ 1. Introduction
While prior research has explored the automation of both physical tasks and digital tasks on user interfaces (Huang et al., 2022; Wang et al., 2023b; Imran et al., 2024; Zhou et al., 2023; Wang et al., 2024; Gur et al., 2023), much of the focus in digital automation has been on enabling AI to understand UI screens, plan specific actions, and navigate through user interfaces (Wen et al., 2023a; Taeb et al., 2023; Wen et al., 2023b; Burns et al., 2022b; Zhang et al., 2021). However, one critical aspect remains underexplored: understanding the consequences of agents performing actions on UIs, especially those that may be risky or irreversible, even when directed by high-level user instructions. In today’s digital age, where interactions with the world are increasingly mediated through screens, it is crucial for AI agents to anticipate and reason about the potential impacts and consequences of their actions on users and their environments.
@@
This paper presents a significant step towards addressing this gap by first constructing a comprehensive taxonomy of UI action impacts on mobile devices, which are central to users’ daily interactions. Through a series of workshops with 12 domain experts in LLMs, mobile UI understanding, and AI safety, we iteratively developed this taxonomy, identifying and categorizing mobile UI actions with potentially significant effects from different domains like impact on the user themselves, impact on other users, the reversibility of this action, and others. In this work, we focus on the impacts caused by attempting to carry out human-initiated directions. In our work, we exclude potential passive impacts, such as those caused by recommendation algorithms that change what users see based on their browsing activities, as such passive impacts are subtle and hard to quantify.
@@
After we established the taxonomy, we conducted a data synthesis study to collect realistic UI actions likely to have impacts by recording UI screen traces and users’ intended actions. We show how existing UI datasets like MoTIF (Burns et al., 2022a) and AndroidControl (Li et al., 2024) contain mostly “harmless” browsing and searching tasks, while our generated UI action traces contain tasks with potential impacts (e.g., changing account information, sending a message).
@@
We collected and labeled 250 mobile UI action traces using our taxonomy via crowdsourcing. Together with 1319 pieces of data sampled from MoTIF and AndroidControl, we evaluated whether LLMs can predict the impact level in comparison to a human’s perceived judgment and accurately classify categories as outlined in our taxonomy. We differentiated the impact levels based on the degree of user intervention required, ranging from users comfortable with an action being automated to those necessitating direct user control. We found that knowledge of our taxonomy could increase an LLM’s ability to predict the overall impact level. However, while existing LLMs demonstrate a moderate understanding on some categories in the taxonomy and predict reasonable results, they still fail to understand all of the nuanced aspects of a mobile UI action’s potential impact. We also present additional findings on how the LLMs we tested overestimated impact and the differences in perceived impact.
@@
  • The development of a comprehensive taxonomy categorizing the impact of mobile UI actions,
  • A collection of human-synthesized mobile UI action traces that include actions with potentially negative impacts, annotated together with data sampled from two other existing datasets, and
  • An evaluation and discussion of how a representative set of current LLMs assess the impact of mobile UI actions, showing a significant gap in their ability to reliably understand the complexity of the impact of mobile UI actions.
@@ 2.1. AI Agents and Safety
LLM-based techniques are moving from limited chat bots to agents (Park et al., 2023; Wang et al., 2023a; Liu et al., 2023b, c; Ahn et al., 2022; Driess et al., 2023a; Li et al., 2023b; Zhang et al., 2024b; Song et al., 2024; Zhang et al., 2024c; Vu et al., 2024) which have the capability to interact with both physical and digital environments. The functionality of these agents also varies based on their initiatives and agencies. Extensive works have explored using LLMs as assistants for creativity tasks including writing (Lee et al., 2022; Chakrabarty et al., 2022), coding (Wu et al., 2024; Agarwal et al., 2024), and content generation (Chang et al., 2023; Lin et al., 2023; Kondratyuk et al., 2023). In the meantime, a number of research efforts have started to explore autonomous AI agents with stronger initiatives and agencies. These agents have been employed in tasks including autonomous decision-making in dynamic environments (Wu et al., 2023; Hong et al., 2023b; Liu et al., 2023d), adaptive problem-solving (Lu et al., 2023; do Nascimento et al., 2023), conducting research (Dai et al., 2023b), manipulating digital interfaces (Wen et al., 2023a; Ruan et al., 2023a), and even interacting with the real world via robotics and IoT systems (Xiao and Wang, 2023; Yang et al., 2023; Cui et al., 2023). These advancements brought us back to the discussion which emerged in earlier years of AI and HCI research, i.e., the dynamics between users and system initiatives (Horvitz, 1999; Hartmann, 2009; Amershi et al., 2019). Fitts first discussed the tradeoff between user initiative and machine agency (Fitts, 1951) by emphasizing the design of systems that align with human cognitive and physical capabilities to improve operator performance and reduce errors. It was followed by the allocation frameworks proposed by Sheridan et al. (Sheridan, 1978) and Parasuraman et al. (Parasuraman et al., 2000), where humans were assigned high-level decision-making and unexpected situation handling tasks, and computers assigned with executing repetitive and precise control actions. Shneiderman (Shneiderman, 2020) also argued more recently that initiative and automation are a two-dimensional construct that humans and AI can both possess high or low levels. Muller and Weisz (Muller and Weisz, 2022) also extended this view by showing how initiative levels can dynamically change within an application. Another related area is robotic process automation (RPA) and its risks in the automation process (Hong et al., 2023a; Eulerich et al., 2024). However, our taxonomy focuses on categorizing the impacts of UI tasks, enabling systems to dynamically shift agency between AI agents and humans depending on the impact and requirements of the task, rather than directly addressing the contexts of where the agency shifts.
More recently, as LLM-based agents’ capabilities have been advanced in different tasks, it is more crucial to address the problem of identifying and reacting to consequences or risks of agent actions, and introduce human confirmation or intervention when necessary. Works focusing on the trustworthiness and safety (Sun et al., 2024; Bai et al., 2022; Glaese et al., 2022) of such agents include ToolEmu (Ruan et al., 2023b) which used an language model emulated sandbox to identify the risks of agents, and TrustAgent (Hua et al., 2024) which used three strategic planning to inject safety knowledge, generate safe plans, and inspect post-planning. Prior work also introduced taxonomies of impact, risk, and potential harms from AI (Cui et al., 2024; Weidinger et al., 2022), but they mostly summarize general risks posed by generative AI, which are often broadly scoped or focused on distinct domains (e.g., content generation, bias amplification). In contrast, our taxonomy focuses uniquely on the real-world impacts of mobile UI actions, as people rely heavily on mobile devices for interacting with the world. It also shares common risks like privacy and security impacts with these taxonomies.
@@ 2.2. UI Understanding, Planning, and Navigation
Prior work has largely explored UI understanding, planning, and navigation on UI screens, which provided a substantial grounding for this work. Prior to the era of LLMs, earlier works in this area studied simpler web and mobile UI screens (Burns et al., 2022a; Gur et al., 2018; Li et al., 2020; Liu et al., 2018; Shi et al., 2017; Wang et al., 2021). Zhang et al. proposed Screen Recognition (Zhang et al., 2021), which predicts bounding boxes, labels, text content, and clickability of UI elements from screenshot pixels. More recently, both LLM (et al., 2023; Driess et al., 2023b; Gu and Dao, 2023; Huang et al., 2023; Jiang et al., 2023a; OpenAI, 2023; Touvron et al., 2023) and multimodal LLM (MLLM) (Dai et al., 2023a; Li et al., 2023c, a; Liu et al., 2023a; McKinzie et al., 2024; Sun et al., 2023; Ye et al., 2023; Zhu et al., 2023) research have advanced the capabilities of UI understanding tasks, including ILuvUI (Jiang et al., 2023b) and Spotlight (Li and Li, 2022) which focus on single-screen UI tasks like screen summarization and widget interaction via GPT-generated data. Ferret-UI (You et al., 2023) introduces strong referring, grounding, and reasoning capabilities to the UI domain.
@@ 3. Building a Taxonomy of UI Impacts
To better understand and categorize the consequences of an autonomous agent taking actions on an app, we created a preliminary taxonomy of the impacts of UI actions. It served as the initial version for later iteration. We then conducted a set of workshop studies to elicit ideas on defining and categorizing UI action effects with 12 participants with expertise in LLMs, AI safety, and UI understanding. Over 4 workshop sessions, participants iterated our initial taxonomy into a refined version.
@@ 3.1. Author-designed Preliminary Taxonomy
Prior to the workshop sessions, the research team met and developed an initial version of a taxonomy designed to categorize the potential impacts of UI actions. The team achieved the initial taxonomy through a pilot workshop study to look at screenshots from existing datasets (AndroidControl (Li et al., 2024) and MoTIF (Burns et al., 2022a)) and brainstorm factors that would affect their decisions on why or why not an action has impact. To make it possible to model and categorize the broad concept of “impact,” the research team grounded the definition of “impact” to be any real world consequences that result directly from active user-initiated actions (e.g., clicking a confirm button or taking a screenshot). We define “action” here as a sequence of singular UI operations, like tapping or swiping, leading to a final execution, like clicking a submission button. Any subtle or passive impact caused by underlying technologies behind the scene is not within our research scope. For example, scrolling through social media might cause future recommended feeds to change, but we are excluding these kind of impacts which are not caused actively by users. The goal of creating our taxonomy was to categorize and understand the consequences of user-initiated UI actions. We structured this initial taxonomy around the following key domains:
@@ 3.2.1. Participants
We recruited 12 domain experts from a large technology company through internal message boards and snowball sampling. The recruitment process began by contacting researchers with expertise in AI, and subsequently extended to those with specialized knowledge in AI safety and UI understanding. A user studies review board internal to our company reviewed and approved our study.
@@ 3.3. Taxonomy Iteration
In this section, we present the findings from our workshop sessions, highlighting the depth and breadth of our discussions on UI action impacts. The taxonomy we developed is intended to cover a wide range of scenarios that extend beyond just determining whether an AI agent should seek human confirmation or intervention. Instead, it broadly considers the various consequences that may arise after a UI action is executed, ensuring its applicability to diverse use cases.
@@ 4.1.2. Apparatus
Figure 4. An annotated example of a monetary transaction. Each of the category is from the taxonomy, with square brackets indicating possible multiple selections of this category.
@@ 4.1.3. Procedure
4.2. Annotation
After collecting the UI action traces, which included both UI action traces and descriptions of the intended actions, we annotated them using our taxonomy through crowdsourcing. The resulting labeled dataset, which includes both synthesized data, and a subset of the AndroidControl and MoTIF datasets, enabled us to evaluate the ability of AI agents to reason about the consequences of UI actions.
Additionally, we asked the annotators to to rate the impact of each action as minimal, moderate, or significant. We carefully defined the different levels of impact in the training process. We asked the annotators to rate an action as minimum impact if they felt like this action can be safely executed automatically without their confirmation. We asked them to rate the action as moderate impact if they felt like they have some concerns over this action and would like to confirm or receive summary of the action before execution. The summary of the action refers to situations when actions contain rich information on the screen that require summarization and confirmation (e.g., shopping cart items before purchasing). Lastly, we asked the annotators to rate the action as significant impact if they felt like they would not allow this action to be executed automatically and would like to intervene themselves. A typical action with significant impact is deleting an account that cannot be restored, where annotators felt that they themselves should be executing it if needed.
Each task was annotated by two annotators. In cases of disagreement, a third annotator was brought in to resolve the discrepancy. The third annotator was responsible for not only providing an additional opinion on the classification and rating questions, but also addressing the cases when the justifications were not similar. While the annotation process was robust, we acknowledge that categorizing interpretive tasks presents significant challenges, even for trained annotators. To assess reliability, we computed the agreement percentages between the initial two annotators across all taxonomy categories. The agreement percentages ranged from 51.6% to 77.9% for multi-label categories like “User Intent” (51.6%), “Impact on UI” (62.3%), “Impact on Self” (69.4%), and “Impact on Others” (77.9%). For single-label categories, the agreement percentages ranged from 55.5% to 98.1%, including “Reversibility” (55.5%), “Roll Back Effects” (90.5%), “Idempotency” (95.8%), “Statefulness” (98.1%), and the general impact level (95.0%). These different levels of agreement reflects the nuanced and interpretive nature of the task. For example, subjective categories like “Impact on Others” and other impact-related categories had lower agreement percentages under 80%. Such complexity underscores the challenges of using this dataset as a benchmark for evaluating large language models (LLMs), and highlights the taxonomy’s value for capturing real-world impacts. By reporting agreement metrics across categories, we aim to provide a transparent foundation for future research and improvements in both annotation processes and model performance.
5. Evaluation: Can LLMs Understand Impact?
In this section, we aim to answer this research question: How do current LLMs understand the impact of UI actions? We summarize the collected data which potentially has impacts and quantitatively compare two existing datasets. We then evaluate five state-of-the-art LLMs with four different LLM prompting strategies to assess their abilities to determine UI action impact based on the taxonomy and to classify specific categories in the taxonomy.
5.1. Data Summary
@@
We collected 250 unique UI action traces from various domains, including e-commerce, social media, financial services, productivity tools, travel and transportation, smart-homes, health, lifestyle, and others. After cleaning identical or highly similar screenshots that were accidentally captured in the collection phase, there are on average 5.3 screens per action trace. In total we collected 1328 UI screens. In the data annotation phase, annotators marked 41 action traces as incomplete, leaving 209 data pieces with annotations. Within the collected data, annotators deemed 23.9% of them to have significant impact and 49.3% of them to have moderate impact, while the remaining 26.8% have minimum impact.